OpenStudy (anonymous):

Needs Checking 5x/2x - x/(x-4) 5x(x-4)/2x(x-4) + -x*2x/2x(x-4) 5x^2-20x/2x(x-4) + -2x^2/2x(x-4) 5x^2-20x+ -2x^2/2x(x-4) 3x^2-20x/2x(x-4)

OpenStudy (anonymous):

@Loser66 Did I make the same mistake as last time?

OpenStudy (loser66):

Yes, you did hahaha ... but at another place. The last line, factor x out and cancel out with the x in denominator to get simpler form

OpenStudy (anonymous):

So \[x(3x-20)\] and the x would make the 2x become just x yes?

OpenStudy (loser66):

\[\dfrac{5\cancel{x}}{2\cancel{x}}-\dfrac{x}{x-4}\\=\dfrac{5(x-4)-2x}{2(x-4)}\\=\dfrac{3x-20}{2(x-4)}\]

OpenStudy (anonymous):

I will get this down yet! Thank you sir :)

OpenStudy (loser66):

I always simplify if I can right after seeing the expression. It makes the process simpler.

OpenStudy (anonymous):

@Loser66 So \[\frac{ 5x }{ 2x }+\frac{ x+1 }{ 2x } = \frac{ 6x+1 }{ 2x } =3+1=4\]

OpenStudy (loser66):

nope, how???

OpenStudy (loser66):

just 6x+1/2x Done.

OpenStudy (loser66):

or \(3+\dfrac{1}{2x}\) You CAN'T break the fraction the way you did. It's wrong